what is the slope of x+2y=8?






















1-what is the slope of a horizontal line?
2-what is the slope of x+2y=8?
thank you









Answer by stanbon(75887)

what is the slope of x+2y=8?
---
2y = -x+8
y = (-1/2)x + 4
------------------------
1-what is the slope of a horizontal line?
Zero
--------------------------
2-what is the slope of x+2y=8?
answered above
===============
Cheers,
Stan H.
==============